The time taken by Arushi to cover 150 km was 2hrs and 30 min more than the time taken in return journey. If she returned at a speed of 10km/hr more than the speed of going,find her speed in each direction.
Answer by stanbon(75887)

The time taken by Arushi to cover 150 km was 2hrs and 30 min more than the time taken in return journey. If she returned at a speed of 10km/hr more than the speed of going,find her speed in each direction.
--------
Going DATA:
dist = 150 km ; rate = r km/hr ; time = 150/r hrs
------------------
Return DATA:
dist = 150 km ; rate = r+10 km/hr ; time = 150/(r+10) hrs
-------
Equation:
time - time = 2.5 hrs
150/r - 150/(r+10) = 2.5 hrs
--------
150r + 150+10 - 150r = 2.5r(r+10)
--------
2.5r^2 + 25r - 1500 = 0
r^2 + 10r - 600 = 0
---
Factor:
(r+30)(r-20) = 0
----
Positive solution:
rate = 20 km/hr (going speed)
r + 10 = 30 km/hr (return speed)
====================================
Cheers,
Stan H.
==============